What is the magnitude and direction of the vector {-5,12}? The direction is measured from the positive x-axis.
Answer by stanbon(75887)

What is the magnitude and direction of the vector {-5,12}? The direction is measured from the positive x-axis.
For any vector (a,b):
magnitude = sqrt(a^2+b^2)
direction = arctan(b/a)
For your problem:
magnitide = sqrt(5^2+12^2)= 13
direction = arctan(-12/5)= 112.62 degrees
Cheers,
Stan H.